A new car is purchased for 24800 dollars. The value of the car depreciates at 12% per year. What will the value of the car be, to the nearest cent, after 13 years?



1
Expert's answer
2022-03-23T02:19:51-0400

According to the compound interest formula

"R=P(1+{\frac i n})^{t*n}" , where P - initial amount, i - annual interest rate, n - number of payments per year, t - number of years, R - amount after t years. So, in the given case we have

"R=24800(1-0.12)^{13}\approx4706.81" dollars will car cost after 13 years